# Topological materials with extensive flat-band surface states

### v2

@@ Abstract
$\frac{1}{2}(1+\sigma_{y})|x_{N}\rangle=\lambda^{N-1}e^{i\phi(N-1)}|x_{1}\rangle$
where $\phi$ is defined by $\Delta_{1}+i\Delta_{2}=\lambda^{i\phi}$. When $\lambda<1$, one solution of this equation for large $N$ is $|x_{N}\rangle\approx 0$ even when $\langle x_{1}|x_{1}\rangle\sim 1$. Hence, Eq. (20) is satisfied, although approximately, with $E=0$, indicating the presence of zero-energy state at the edge $|x_{1}\rangle$.
Critical lines As mentioned in Fig. 1 of the main text, the non-analyticity at the phase boundaries is in the form of a discontinuity in $\partial^{3}E_{0}/\partial t^{3}$ in some cases and in the form of divergence of $\partial^{3}E_{0}/\partial t^{3}$ in the other cases. The amount of discontinuity varies along the boundary line. The line $J=1$ forms the boundary (II, V) and the one (III, IV). We show the variation of the amount of discontinuity along this line in Fig. (10(a)). Similarly, the line $J-2t=-1$ forms the boundary (II, III) and the one (IV, V). The variation of the amount of discontinuity along this line is shown in Fig. (10(b)). On the other hand, the third derivative $\partial^{3}E_{0}/\partial t^{3}$ diverges algebraically along the boundaries (I, II) and (V, VI), with the exponent depending upon the direction of approach. This is shown in Figs. (11, 12).
@@
Figure 11: Algebraic divergence of $\partial^{3}E_{0}/\partial J^{3}$ at the critical line $J-2t=1$ : The dots are the numerical data and the green line is the fitted curve of the form $ax^{b}$. The critical exponent is $b=0.11$ when the critical line is approached from below and $b=0.14$ when the critical line is approached from above. $t$ is fixed at $0.2$. Both $x$ and $y$ axes are in log scale.
Figure 12: Algebraic divergence of $\partial^{3}E_{0}/\partial J^{3}$ at the critical line $J+2t=1$ : The dots are the numerical data and the green line is the fitted curve of the form $ax^{b}$. The critical exponent is $b=0.19$ when the critical line is approached from below and $b=0.22$ when the critical line is approached from above. $t$ is fixed at $0.3$. Both $x$ and $y$ axes are in log scale.
Dispersion relations of Phase I with disorder We find that a random fluctuation of the hopping amplitude around 1.0 by 10% does not destroy the zero energy surface states (Fig. 13).
@@
Figure 13: Dispersion relation for Phase I : The dispersion relations are calculated for $J=0.2$ and $t=0.25$. The disorder is put in the Z direction where it fluctuates uniformly and randomly between $0.9$ and $1.1$.

### v3

@@ Abstract
$\frac{1}{2}(1+\sigma_{y})|x_{L_{z}}\rangle=\lambda^{{L_{z}}-1}e^{i\phi({L_{z}}-1)}|x_{1}\rangle$ (22)
where $\phi$ is defined by $\Delta_{1}+i\Delta_{2}=\lambda e^{i\phi}$. When $\lambda<1$, one solution of this equation for large $L_{z}$ is $|x_{L_{z}}\rangle\approx 0$ even when $\langle x_{1}|x_{1}\rangle\sim 1$. Hence, Eq. (19) is satisfied, although approximately, with $E=0$, indicating the presence of zero-energy state at the edge $|x_{1}\rangle$. These states will not be present when $\lambda>1$. Thus, the condition for the existence of zero-energy states is
$\lambda<1,\;\;\;{\mbox{o}r}\;\;\;\Delta_{1}^{2}+\Delta_{2}^{2}<1$ (23)
Analogy with SSH model: As mentioned above, the topological aspect of our model has resemblance with Su-Schrieffer- Heeger model [SSH]. To be more specific, we note that the Hamiltonian for this model can be written as
${\mathcal{H}}_{k}=(\Gamma+\cos k)\sigma_{x}+\sin k\,\sigma_{y}$
where $k$ is the (one-dimensional) wave-vector and $\Gamma$ is the ratio of hopping amplitudes. The energy
$E(k)=\sqrt{\Gamma^{2}+1+2\Gamma\cos k}$ (24)
shows a singularity at the point $\Gamma=1$, where there is a transition between topologically trivial and non-trivial phases. In our case, the energy in Eq. (10)
$E(\vec{k})\equiv\sqrt{A_{\vec{k}}^{2}+B_{\vec{k}}^{2}}=\sqrt{(\Delta_{1}-\cos k_{z})^{2}+(\Delta_{2}+\sin k_{z})^{2}}$ (25)
can be written as
$E(\vec{k})=\sqrt{\lambda^{2}+1+2\lambda\cos(k_{z}-\phi)}$ (26)
The similarity between Eqs. (24) and (26) indicates that the topological transition and singularity at $\lambda=1$
Critical lines: As mentioned in Fig. 1 of the main text, the non-analyticity at the phase boundaries is in the form of a discontinuity in $\partial^{3}E_{0}/\partial t^{3}$ in some cases and in the form of divergence of $\partial^{3}E_{0}/\partial t^{3}$ in the other cases. The amount of discontinuity varies along the boundary line. The line $J=1$ forms the boundary (II, V) and the one (III, IV). We show the variation of the amount of discontinuity along this line in Fig. (10(a)). Similarly, the line J2t=1J-2t=-1 forms the boundary (II, III) and the one (IV, V). The variation of the amount of discontinuity along this line is shown in Fig. (10(b)). On the other hand, the third derivative $\partial^{3}E_{0}/\partial t^{3}$ diverges algebraically along the boundaries (I, II) described by equation $J+2t=1$ and (V, VI) described by equation J2t=1J-2t=1, with the exponent depending upon the direction of approach. This is shown in Figs. (11, 12). It needs to be mentioned that non-analyticities arise in the third derivative of the ground state energy and first and second derivatives does not have non-analyticities. So these phase transitions are neither first nor second order phase transitions [Wu_2023].
@@
Figure 11: Algebraic divergence of $\partial^{3}E_{0}/\partial J^{3}$ at the critical line J2t=1J-2t=1 separating V and VI : The dots are the numerical data and the dashed line is the fitted curve of the form $ax^{-b}$. The critical exponent is $b=0.11$ when the critical line is approached from below and $b=0.14$ when the critical line is approached from above. $t$ is fixed at $0.2$. Both $x$ and $y$ axes are in log scale.
Figure 12: Algebraic divergence of $\partial^{3}E_{0}/\partial J^{3}$ at the critical line $J+2t=1$ separating phase I and II : The dots are the numerical data and the dashed line is the fitted curve of the form $ax^{-b}$. The critical exponent is $b=0.19$ when the critical line is approached from below and $b=0.22$ when the critical line is approached from above. $t$ is fixed at $0.3$. Both $x$ and $y$ axes are in log scale.
Dispersion relations of Phase I with disorder: We find that a random fluctuation of the hopping amplitude around 1.0 by 10% does not destroy the zero energy surface states (Fig. 13).
@@
Figure 13: Dispersion relation for Phase I : The dispersion relations are calculated for $J=0.2$ and $t=0.25$. The disorder is put in the Z direction, where it fluctuates uniformly and randomly between $0.9$ and $1.1$.
Bulk dispersion relation of Phase I: Here we present the dispersion relation for the bulk Hamiltonian in Phase I($J=0.2$ and $t=0.2$) in Fig. [14].
(a) $k_{x}=k_{y}=0$ and $k_{y}=k_{z}=0$
(b) $k_{x}=k_{y}={\pi\over 2}$ and $k_{y}=k_{z}={\pi\over 2}$
(c) $k_{x}=k_{y}=\pi$ and $k_{y}=k_{z}=\pi$
Figure 14: Dispersion relation for Phase I of the Bulk Hamiltonian Eq (1)